# Generalized saddle-node ghosts and their composite structures in dynamical systems

### 09 Apr 2026 v2

@@ 4 Methods
All models in this study were implemented in Python and simulations were performed using a RK45 integration scheme as implemented in SciPy’s solve_\_ivp function [71]. Numerical bifurcation analysis of the coupled theta neurons and tipping element models was performed using XPPAUT. A detailed description of the GhostID algorithm for identifying ghosts from trajectories is given in subsection 2.2 and Figure 2. Algorithms for phase space sampling, continuing identified ghosts across parameters and identifying composite ghost structures such as ghost channels and ghost cycles are described in Supplementary section 1.2.

### 16 Apr 2026 v3

@@ 4 Methods
All models in this study were implemented in Python and simulations were performed using the RK45 integration scheme as implemented in SciPy’s solve_\_ivp function [71]. Numerical bifurcation analysis of the coupled theta neurons and tipping element models was performed using XPPAUT. A detailed description of the GhostID algorithm for identifying ghosts from trajectories is given in subsection 2.2 and Supplementary Figure 2. Algorithms for phase space sampling, continuing identified ghosts across parameters and identifying composite ghost structures such as ghost channels and ghost cycles are described in Supplementary section 1.2.